OpenStudy (anonymous):

f(x) = 1/2x^2 +5/2x−3/2 (a) Find all the real zeros of the polynomial function. x = (smaller value) x = (larger value)

OpenStudy (anonymous):

am i supposed to start off by isolating x^2?

OpenStudy (phi):

looks like you have to use the quadratic formula set f(x)= 0 1/2x^2 +5/2x−3/2 = 0 (I assume this means (1/2) * x^2 (and not 1 over (2x^2)) multiply both sides of the equation by 2 (this simplifies it a little) x^2 + 5x -3 =0 It does not factor nicely, so we use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \] where a,b,c match the coefficients: \[ ax^2 + bx +c =0\]

OpenStudy (anonymous):

i have a question. if this was a simple (x+2)^2 type of thing, i wouldn't use that formula. only use that formula when it doesn't factor nicely?

OpenStudy (phi):

If you can factor it easily. But the formula works: (x+2)^2=0 is x^2 +4x +4 = 0 a=1, b=4, c=4 so (-4 ±sqrt(16-4*1*4))/2 = (-4 ±0)/2 = -2 so x= -2 or x+2= 0 here the two solutions are both -2

OpenStudy (anonymous):

i got -5+-√37)/2

OpenStudy (anonymous):

thanks for your help!

OpenStudy (phi):

yes , but don't forget the parens (-5+-√37)/2 (2 divides both the -5 and the root)